OpenStudy (anonymous):

Kevin and some college friends are hiking on a trail that is 70 feet above sea level. They hike into a canyon that is 14 feet below sea level. How many feet have Kevin and his friends descended? __ feet

OpenStudy (anonymous):

70-(-14)=70+14=84